OpenStudy (anonymous):

Word Problem PLEASE HELP): WILL GIVE MEDAL! A bungee jumper falls from a bridge 400 feet above ground. Her height above ground after t seconds is given by the equation h=-16t^2+400 where h is the height of the jumper in feet and t is the time in seconds since her jump. Approximately how long will it take for the bungee jumper to reach 100 feet above the ground on her initial jump? A. 14 seconds B. 4.3 seconds C. 4.4 seconds D. 4.5 seconds

OpenStudy (anonymous):

A fireworks display is launched from a platform 10 feet above ground with an initial upward velocity of 70 feet per second. The height of the fireworks above ground after t seconds is given by the equation h = –16t^2 + 70t + 10, where h is the height of the fireworks in feet and t is the time in seconds after they are launched. What is the maximum height of the fireworks display, to the nearest foot?

OpenStudy (anonymous):

A. 87 feet B. 174 feet C. 70 feet D. 10 feet

OpenStudy (anonymous):

answer for first one is b.4.3 h= -16t^2+400 100= -16t^2+400 100-400=-16t^2 -300 = 16t^2 (divide both sides by -16) 18,75 = t^2 (take square root of both sides 4.3 = t

OpenStudy (anonymous):

thank you! by any chance could you help me with one more similar to it? @amyvincent12

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A model jet is fired up in the air from a 16-foot platform with an initial upward velocity of 52 feet per second. The height of the jet above ground after t seconds is given by the equation h=-16t^2+52t+16, where h is the height of the jet in feet and t is the time in seconds since it is launched. Approximately how long does it take the jet to reach its highest point? A. 0.3 seconds B. 1.6 seconds C. 3.3 seconds D. 6.5 seconds @amyvincent12

OpenStudy (anonymous):

do you have a list of equations to use for questions like these? you would need an equation that has.. v1, v2, t and d… ?